OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER!! Solve for w 2(2w+9)=50

OpenStudy (anonymous):

w is 8

OpenStudy (anonymous):

2x8+9 is 25 then 25x2 is 50

OpenStudy (xmissalycatx):

Try to distribute the 2 to everything INSIDE of the parentheses and it'll give you: 4x + 18 = 50 Now you need to SUBTRACT 18 from both sides: 4x + 18 = 50 -18 -18 ______________ 4x = 32 Now divide 4 from both sides: 4x = 32 --- --- 4 4 so x = __?__

OpenStudy (xmissalycatx):

@Ballerina_MJ Do not give direct answers, thank you..

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

Thank you!!!!! :D

OpenStudy (damonte212):

..